The median annual salary for a police dispatcher in a large city was $42,528. If the 25th percentile for the annual salary of a police dispatcher was $34,961, find the percent of police dispatchers with the following annual salaries. (See Example 2 in this section.)

(a)less than $42,528 %

(b)more than $34,961 %

(c)between $34,961 and $42,528 %
Answer #1

Median divide the data into two equal proportions.

a) Percent of police dispatchers with annual salaries less than $42528 = 50%

b) Percent of police dispatchers with annual salaries more than $34961 =100-25 = 75%

c) Percent of police dispatchers with annual salaries between $34961 and $42528 = 50-25 = 25%